# Repetitive Penrose process in Kerr-de Sitter black holes

### v1

@@ Abstract
Recently, references 8 ; 4 found that the repetitive Penrose process cannot extract all the extractable rotational energy of a Kerr black hole, and reference 5 found that the repetitive electric Penrose process cannot extract all the electrical energy of a Reissner-Nordström (RN) black hole. This suggests that a law analogous to the third law of thermodynamics exists for the repetitive Penrose process. In this paper, we intend to study the repetitive Penrose process in the Kerr-de Sitter (Kerr-dS) black hole. We will explore influences of the cosmological parameter on the repetitive Penrose process. The results show that, in addition to a similar third law of thermodynamics, the Kerr-dS black hole yields a higher energy return on investment (EROI) and single-extraction energy capability compared to the Kerr black hole. Specifically, the larger the cosmological parameter, the stronger the EROI and the single-extraction energy capability. Furthermore, we also find that at a lower decay radius, the Kerr black hole exhibits a higher energy utilization efficiency (EUE) and more extracted energy after the repetitive Penrose process is completed. However, at a higher decay radius, the situation is reversed, i.e., the Kerr-dS black hole exhibits a higher EUE and more extracted energy, which is due to the existence of stopping condition of the iteration.
@@ I Introduction
Extracting energy from black holes has long been a topic of great interest to physicists. In 1969, Penrose proposed a method for extracting energy from a rotating black hole 9 . A particle with mass $\mu_{0}$ and energy $E_{0}$ falls from infinity and decays within the ergosphere into two particles, $\mu_{1}$ and $\mu_{2}$. The particle $\mu_{1}$ with negative energy $E_{1}$ falls into the event horizon of the black hole, while the particle $\mu_{2}$ escapes to infinity carrying more energy $E_{2}$ than the original particle. Thus, rotational energy $E_{2}-E_{0}$ is extracted from the black hole. However, Bardeen, Wald, and others pointed out that the original Penrose process lacks a feasible initiation mechanism. In order to extract energy from the black holes, the velocity of the outgoing particle should exceed half the speed of light 10 ; 11 . Building on Penrose’s pioneering work, scientists have discovered many other energy extraction mechanisms, such as the magnetic Penrose process 12 ; 13 , the collisional Penrose process 14 , superradiant scattering 15 , the Blandford–Znajek mechanism 16 , the magnetohydrodynamic Penrose process 17 , the Bañados–Silk–West mechanism 18 , and the Comisso–Asenjo mechanism 19 . In particular, based on the Comisso–Asenjo process, the study of magnetic reconnection mechanisms, as well as the power and efficiency of energy extraction via such mechanisms, have been investigated extensively6 ; 7 ; 26 ; 27 ; 28 ; 29 ; 30 .
@@
On the other hand, Misner, Thorne, and Wheeler proposed a repetitive Penrose process 20 . They envisioned an advanced civilization constructing a rigid framework around a black hole and building a vast city on it. Daily, trucks would transport one million tons of garbage from the city to a dumping ground. At the dump, the garbage would be loaded into shuttles, which would then be dropped one after another toward the center of the black hole. In terms of the Penrose process, this means repeating the decay in the same manner, one after another, into the black hole. However, Ruffini et al. 8 found very recently that such a process would violate energy conservation. This is because the effect of the irreducible mass 21 ; 2 ; 3 was not taken into account in the process. For each iteration, the mass and spin of the black hole were not updated, treating the iterative process as a linear one. In light of this, references 8 ; 4 proposed a modified repetitive Penrose process, wherein for each iteration, the updated mass and spin of the black hole are used to extract energy, while also accounting for the new irreducible mass. This renders the iterative process nonlinear. They discovered that in the repetitive Penrose process, it is impossible to extract all the extractable energy of a Kerr black hole, and the reduction in extractable energy is primarily converted into irreducible mass. Currently, the work of references 8 ; 4 has been extended to the electric Penrose process 5 , where it was found that the repetitive electric Penrose process cannot extract all the electrical energy of a RN black hole, indicating a like-third law of thermodynamics holds.
@@
Inspired by the works in references 8 ; 4 , this paper will investigate the repetitive Penrose process in Kerr-dS spacetime. Kerr-dS spacetime is an exact solution to the Einstein field equations with a positive cosmological constant. It describes a rotating black hole embedded in a de Sitter universe. Currently, a growing body of research focuses on a positive cosmological constant, such as cosmic inflation 22 , phase transitions 23 , thermodynamics 24 , dS/CFT correspondence 25 , and so forth. In Kerr-dS spacetime, besides verifying the existence of a like-third law of thermodynamics, we will also explore the influence of the cosmological parameter on the repetitive Penrose process. In particular, we will compare our results with those in the Kerr black hole in terms of the EROI, single-extraction energy capability, EUE, and the variation in total extracted energy. The results indicate that, in addition to a like-thermodynamic third law, Kerr-dS spacetime possesses a a higher EROI and single-extraction energy capability compared to Kerr spacetime. The larger the cosmological parameter, the stronger the EROI and single-extraction energy capability. In the case of a lower decay radius, after the completion of the repetitive Penrose process, the Kerr black hole exhibits a higher EUE and yields more extracted energy. However, in the case of a higher decay radius, the situation is reversed. That is, the Kerr-dS black hole exhibits a higher EUE and yields more extracted energy.
@@ II The Penrose Process in Kerr-dS Black Holes
This inequality constrains the choices of the cosmological parameter and the spin. In this paper, to maximize energy extraction, we will consider the repetitive Penrose process for extremal Kerr-dS black holes on the equatorial plane. That is, for a given cosmological parameter, the initial spin corresponds to the extremal black hole case, i.e., satisfying $r_{+}=r_{-}$. Here, $\Lambda M^{2}$ can range from 0 to 0.099555, and $a/M$ ranges from 1 to 1.04075.
The irreducible mass of the black hole can be determined by the formula for the area of the black hole’s event horizon, namely 2 ; 3
$M_{irr}=\sqrt{\frac{S}{16\pi}},$ (5)
where
$S=\int_{0}^{\pi}\int_{0}^{2\pi}\sqrt{g_{\theta\theta}g_{\phi\phi}}\,d\phi d\theta=\frac{4\pi(r_{+}^{2}+a^{2})}{1+\frac{1}{3}\Lambda a^{2}}.$ (6)
Then the extractable energy can be expressed as
$E_{extractable}=M-M_{irr}=M-\sqrt{\frac{r_{+}^{2}+a^{2}}{4\left(1+\frac{\Lambda}{3}a^{2}\right)}},$ (7)
@@
Similarly, according to the horizon equation and the irreducible mass formula, $r_{+}$ and $M_{irr}$ will also change. The change in extractable energy is
$\Delta E_{extractable,n-1}=\Delta M_{n-1}-\Delta M_{irr,n-1}.$ (22)
During this process, the extracted energy is
@@ IV Conclusion
The results show that the repetitive Penrose process cannot extract all the extractable energy from Kerr-dS black holes, indicating the existence of a law analogous to the third law of thermodynamics. Only a small portion of the reduction in extractable energy is converted into extracted energy, while the majority is transformed into irreducible mass. As the iteration is ended, a significant amount of extractable energy remains, and it may be extracted through other means, which aligns with previous conclusions. Furthermore, the Kerr-dS spacetime exhibits a higher EROI and a greater single-extraction energy capability compared to the Kerr spacetime. The larger the cosmological parameter, the stronger the EROI and single-extraction energy capability. In cases with a lower decay radius, after the completion of the repetitive Penrose process, the Kerr black hole demonstrates a higher EUE and yields more extracted energy. However, in cases with a higher decay radius, the situation is reversed, the Kerr-dS black hole shows a higher EUE and yields more extracted energy. This is due to the existence of an iteration stopping condition.

### v2

@@ Abstract
Recently, references 8 ; 4 found that the repetitive Penrose process cannot extract all the extractable rotational energy of a Kerr black hole, and reference 5 found that the repetitive electric Penrose process cannot extract all the electrical energy of a Reissner-Nordström (RN) black hole. In this paper, we intend to study the repetitive Penrose process in the Kerr-de Sitter (Kerr-dS) black hole. We will explore influences of the cosmological parameter on the repetitive Penrose process. The results show that, apart from the comparable features documented earlier, the Kerr-dS black hole yields a higher energy return on investment (EROI) and single-extraction energy capability compared to the Kerr black hole. Specifically, the larger the cosmological parameter, the stronger the EROI and the single-extraction energy capability. Furthermore, we also find that at a lower decay radius, the Kerr black hole exhibits a higher energy utilization efficiency (EUE) and more extracted energy after the repetitive Penrose process is completed. However, at a higher decay radius, the situation is reversed, i.e., the Kerr-dS black hole exhibits a higher EUE and more extracted energy, which is due to the existence of stopping condition of the iteration.
@@ I Introduction
Extracting energy from black holes has long been a topic of great interest to physicists. In 1969, Penrose proposed a method for extracting energy from a rotating black hole 9 . A particle with mass $\mu_{0}$ and energy $E_{0}$ falls from infinity and decays within the ergosphere into two particles, $\mu_{1}$ and $\mu_{2}$. The particle $\mu_{1}$ with negative energy $E_{1}$ falls into the event horizon of the black hole, while the particle $\mu_{2}$ escapes to infinity carrying more energy $E_{2}$ than the original particle. Thus, rotational energy $E_{2}-E_{0}$ is extracted from the black hole. However, Bardeen, Wald, and others pointed out that the original Penrose process lacks a feasible initiation mechanism. In order to extract energy from the black holes, the velocity of the outgoing particle should exceed half the speed of light 10 ; 11 . Building on Penrose’s pioneering work, scientists have discovered many other energy extraction mechanisms, such as the magnetic Penrose process 12 ; 13 , the collisional Penrose process 14 , superradiant scattering 15 , the Blandford–Znajek mechanism 16 , the magnetohydrodynamic Penrose process 17 , the Bañados–Silk–West mechanism 18 , the possible energy-extraction Penrose-like mechanism which violates the null energy condition32 ; 33 , and the Comisso–Asenjo mechanism 19 . In particular, based on the Comisso–Asenjo process, the study of magnetic reconnection mechanisms, as well as the power and efficiency of energy extraction via such mechanisms, have been investigated extensively6 ; 7 ; 26 ; 27 ; 28 ; 29 ; 30 ; 31 .
@@
On the other hand, Misner, Thorne, and Wheeler proposed a repetitive Penrose process 20 . They envisioned an advanced civilization constructing a rigid framework around a black hole and building a vast city on it. Daily, trucks would transport one million tons of garbage from the city to a dumping ground. At the dump, the garbage would be loaded into shuttles, which would then be dropped one after another toward the center of the black hole. In terms of the Penrose process, this means repeating the decay in the same manner, one after another, into the black hole. However, Ruffini et al. 8 found very recently that such a process would violate energy conservation. This is because the effect of the irreducible mass 21 ; 2 ; 3 was not taken into account in the process. For each iteration, the mass and spin of the black hole were not updated, treating the iterative process as a linear one. In light of this, references 8 ; 4 proposed a modified repetitive Penrose process, wherein for each iteration, the updated mass and spin of the black hole are used to extract energy, while also accounting for the new irreducible mass. This renders the iterative process nonlinear. They discovered that in the repetitive Penrose process, it is impossible to extract all the extractable energy of a Kerr black hole, and the reduction in extractable energy is primarily converted into irreducible mass. Currently, the work of references 8 ; 4 has been extended to the electric Penrose process 5 , where it was found that the repetitive electric Penrose process cannot extract all the electrical energy of a RN black hole.
@@
Inspired by the works in references 8 ; 4 , this paper will investigate the repetitive Penrose process in Kerr-dS spacetime. Kerr-dS spacetime is an exact solution to the Einstein field equations with a positive cosmological constant. It describes a rotating black hole embedded in a de Sitter universe. Currently, a growing body of research focuses on a positive cosmological constant, such as cosmic inflation 22 , phase transitions 23 , thermodynamics 24 , dS/CFT correspondence 25 , and so forth. In Kerr-dS spacetime, in addition to the impossibility of reaching a state of zero extractable energy through the repetitive Penrose process, we will also explore the influence of the cosmological parameter on the repetitive Penrose process. In particular, we will compare our results with those in the Kerr black hole in terms of the EROI, single-extraction energy capability, EUE, and the variation in total extracted energy. The results indicate that, in addition to exhibiting previously observed similar phenomena, Kerr-dS spacetime possesses a higher EROI and single-extraction energy capability compared to Kerr spacetime. The larger the cosmological parameter, the stronger the EROI and single-extraction energy capability. In the case of a lower decay radius, after the completion of the repetitive Penrose process, the Kerr black hole exhibits a higher EUE and yields more extracted energy. However, in the case of a higher decay radius, the situation is reversed. That is, the Kerr-dS black hole exhibits a higher EUE and yields more extracted energy.
@@ II The Penrose Process in Kerr-dS Black Holes
This inequality constrains the choices of the cosmological parameter and the spin. In this paper, to maximize energy extraction, we will consider the repetitive Penrose process for extremal Kerr-dS black holes on the equatorial plane. That is, for a given cosmological parameter, the initial spin corresponds to the extremal black hole case, i.e., satisfying $r_{+}=r_{-}$ [^2] What needs to be emphasized here is that by ”extremal black hole,” we refer to the initial black hole being extremal, and not that it remains extremal after iterations of the repetitive Penrose process. This is consistent with the approach in reference 4 , where they consider an initial extremal Kerr black hole with $a=M$, which does not remain extremal after subsequent iterations.. Here, $\Lambda M^{2}$ can range from 0 to 0.099555, and $a/M$ ranges from 1 to 1.04075.
Within the event horizon, in addition to the mass $M$ itself, there exists a local positive energy contribution due to the cosmological constant 34 . A local, continuous, positive-definite effective mass function can be expressed as 34
$M_{loc}(r)=M+\frac{\Lambda r}{6}\left(r^{2}+a^{2}\right).$ (5)
Thus, the irreducible mass can be written as 34
$M_{irr}=\sqrt{\frac{1}{2}M_{loc,H}(M_{loc,H}+\sqrt{M_{loc,H}^{2}-a^{2}})}.$ (6)
Here, $M_{loc,H}$ denotes the value of $M_{loc}(r)$ at the event horizon. Then the extractable energy can be expressed as
$E_{extractable}=M_{loc,H}-M_{irr},$ (7)
@@
Similarly, according to the horizon equation, the irreducible mass formula, and the expression for extractable energy, $r_{+},M_{irr}$ and $E_{extractable}$ will also change. During this process, the extracted energy is
@@ IV Conclusion
The results show that the repetitive Penrose process cannot extract all the extractable energy from Kerr-dS black holes. Only a small portion of the reduction in extractable energy is converted into extracted energy, while the majority is transformed into irreducible mass. As the iteration is ended, a significant amount of extractable energy remains, and it may be extracted through other means, which aligns with previous conclusions. Furthermore, the Kerr-dS spacetime exhibits a higher EROI and a greater single-extraction energy capability compared to the Kerr spacetime. The larger the cosmological parameter, the stronger the EROI and single-extraction energy capability. In cases with a lower decay radius, after the completion of the repetitive Penrose process, the Kerr black hole demonstrates a higher EUE and yields more extracted energy. However, in cases with a higher decay radius, the situation is reversed, the Kerr-dS black hole shows a higher EUE and yields more extracted energy. This is due to the existence of an iteration stopping condition.